### v1

@@ 3.3 Evaluation Setup
All military test sets produce freeform outputs that are graded with a judge model. In this work we used the W8A8 quantized version of the Atla Selene 1 model Alexandru et al. (2025), which is based on Llama 3.3 70B Grattafiori et al. (2024) and is further trained using Supervised Fine Tuning (SFT) and Direct Preference Optimization (DPO) to be a better judge. To remove additional variables, we use Inspect’s default prompt for the LLM judge, and we do not use any system or developer prompt for any of the military tasks. Further details on the evaluation setup are provided in Appendix B.
@@ 4.3 Hyperparameter Analysis
    Chat Template: Most of our experimentation was performed using the gpt-oss-20b chat template. However, we also tried using a substantially different chat template for training, the alpaca template (see Appendix A). Inference was still conducted using the original gpt-oss-20b chat template. We found that using the alpaca chat template for training provided a small but significant boost in performance, again due to less degradation of the reasoning behavior of the model.
@@ 6 Conclusion and Next Steps
Appendix A Training Setup
@@
Appendix B Evaluation Setup

### v2

@@ 3.3 Evaluation Setup
All military test sets produce freeform outputs that are graded with a judge model. In this work we used the W8A8 quantized version of the Atla Selene 1 model Alexandru et al. (2025), which is based on Llama 3.3 70B Grattafiori et al. (2024) and is further trained using Supervised Fine Tuning (SFT) and Direct Preference Optimization (DPO) to be a better judge. To remove additional variables, we use Inspect’s default prompt for the LLM judge, and we do not use any system or developer prompt for any of the military tasks. Further details on the evaluation setup are provided in Appendix C.
@@ 4.3 Hyperparameter Analysis
    Chat Template: Most of our experimentation was performed using the gpt-oss-20b chat template. However, we also tried using a substantially different chat template for training, the alpaca template (see Appendix B). Inference was still conducted using the original gpt-oss-20b chat template. We found that using the alpaca chat template for training provided a small but significant boost in performance, again due to less degradation of the reasoning behavior of the model.
@@ 6 Conclusion and Next Steps
Appendix A Test Set Examples
Examples from our four test sets are provided in Table A.
Table 9: Samples from each of our four tests.
[1pt] Test Set Question Answer
[0.5pt] combat-arms How do artillery units provide counterfire support in a contested EW environment? By using preplanned fire missions, manual gunnery solutions, redundant comms, and passive sensors to detect enemy artillery.
combat-medic According to ATP 4-25.12 (Unit Field Sanitation Team), what are the steps a 68W Combat Medic should take to ensure proper disease prevention measures in a deployed environment? Under ATP 4-25.12, a 68W Combat Medic should assess the operational environment, implement unit-level field sanitation measures for disease prevention measures, coordinate with the Unit Field Sanitation Team, and provide preventive medicine guidance to reduce disease and non-battle injuries, ensuring soldier readiness.
cyber Explain how ethical hacking engagements strengthen organizational resilience. Ethical hacking involves authorized testing to identify weaknesses before adversaries can exploit them. Results inform patch management, training, and policy revisions. By exploiting gaps safely, ethical hackers validate incident response capabilities and measure defense depth. Regular testing fosters a security culture and continuous improvement loop.
mil-bench-5k How does the concept of “systemic disruption” contribute to defeating adversaries in urban operations, and what are the key mechanisms involved? Systemic disruption serves as a decisive strategy in urban operations by degrading an adversary’s capacity to function as an integrated force within complex urban environments. This approach prioritizes: 1. **Disruption of Support Systems** - Isolating adversaries from critical resources (supplies, communication networks, financial infrastructure) - Severing ties to local support networks through population engagement and counterintelligence 2. **Precision Force Application** - Targeted use of kinetic and non-kinetic effects to degrade adversary capabilities while minimizing infrastructure damage - Leveraging advanced ISR (Intelligence, Surveillance, Reconnaissance) for high-value target identification 3. **Population-Centric Operations** - Building trust with local populations to deny adversary recruitment and logistical access - Delivering essential services to establish legitimacy and counter adversary narratives By combining these mechanisms, joint forces can collapse adversarial influence while preserving urban systems’ functionality, aligning with the principles of proportionality and military necessity in densely populated areas.
[1pt]
Appendix B Training Setup
@@
Appendix C Evaluation Setup